step by step pleasefreda ſ sin sin(x) dx. 3.r? – 112 + 10 dr. 6.2? + 10.2 < V3 - 2.0 - x? dr. In(a) dr if it exists. .c?
Answer #1

. dn. dn - 6n+10 n6n +con +3° -3° +10 n - 60+10 x - 2.3.n. 2 n-Gatio = dn. 1-3) +1 Let UE n-3 du= dr. S v²t ti Tant (u) Tantan Cos (22) Constant einen) +20 +20 T Constat sin (2n) + 2a + sinn. Coen 4an ein (u) 11 a sin (sin o) Cas (sin lo) + 2 sint()(ا = - = Ina ,ه د (( د ) ع ) سے۔ - 1) Ina Cim ē. (uri) (9 -Inz ((+1) ك ب 0 ا+ =f) (cim (ut) (2 + 1 مل). ( را Lindeterminate fè da. sfon. gen) dx= fx) Sganda - Sa tw. Samide.de formula. fem) = 2 q (a) e Here 11 had - on. Seda - Soa). sede de - Sax. dx= p ) ( - Goen) – 14 4 - Cac (3) + Conetant, 3 - 3 Cosn Cos (3n) Constant 11 + 12 Cos (32) Constant I sin n = -3 Coen + 12 3du 3n 12nt to dn S x3 - 6x2tion In (ult - Constant. se content in 2-cm+10ml + Constad --